OpenStudy (anonymous):

HI i was doing some integration and come across to this. a^4 sin^2(x) cosx cosx = a^4/ 4 (sin^2 (2x))??

OpenStudy (anonymous):

\[a ^{4}\times \sin ^{2}x \cos x \cos x = a ^{4}\div4\times \sin ^{2}2x\]

OpenStudy (anonymous):

i dono how u convert and get this, please help me, tomorrow there is exam

OpenStudy (anonymous):

are you integrating with respect to \(a\) ?

OpenStudy (anonymous):

no, this has nothing to do with integration yet, this is simplying the equation to integrate it, but i don understand how they get the end answer

OpenStudy (anonymous):

because u can integrate the right hand side, just by changing sin^2x to 1/2 - 1/2 cosx, but i don understand how they get the sin^2x from the left hand side

OpenStudy (anonymous):

nvm solved (: